A 0.5-g sample of TNT(C7H5N2O6) is burned in a bomb calorimeter containing 610 g of water at an initial temperature of 20°C. The heat capacity of the bomb is 420 J/°C and the heat of combustion of TNT is 3374 kJ/mol. Calculate the final temperature of the water and calorimeter once the reaction is complete.


1
Expert's answer
2021-08-30T02:25:49-0400

Heat lost = Heat gained

"n_t \u00d7 H_c = C_b\u2206T_b+ m_wc_w\u2206T_w"


number of moles of TNT = 0.5/213.12 = 0.00235


"0.00235× 3374000 = 420× (T_f-20) + 0.610 × 4200× (T_f- 20)\\
7928.9= 420T_f - 8400 + 2562
T_f - 51240
\\
 2982T_f= 67568.9\\
T_f= 22.67°C"